Circular No. 6247

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVAE IN 1989 C. Pollas, Observatoire de la Côte d’Azur (OCA), reports his discov-

ery of two apparent supernovae on photos taken in the course of the second Palomar Sky Survey (PSS). SN 1989ad is located at α = 13^{h}25^{m}38^{s}.96, δ = +32^{o}34′51′′.7 (equinox 2000.0), or 2′′.9 east and 11′′.5 north of the host galaxy’s center. Estimated magnitudes for SN 1989ad from various photographs: 1989 Feb. 7.2 UT, B = 19.6 (OCA Schmidt); Feb. 8.2 and 9.2, R = 19.9 (OCA); Mar. 9.8, R = 20.4 (PSS field 383); Mar. 11.1, B = 21.2 (OCA); Mar. 12.1, V = 20.6 (OCA). This star is seen neither on a B PSS film taken in 1991 Apr. nor on OCA plates and films taken in 1990 and 1995. A nearby star (B ∼ 15.5) has position end figures 44^{s}.25, 35′16′′.2. SN 1989ae is located 4′′.0 west and 0′′.5 south of the center of UGC 7122 (α = 2^{h}14^{m}40^{s}.7, δ = +43^{o}21′28′′). Estimated magnitudes for SN 1989ae from the PSS photos: 1989 Aug. 31.93 UT, B ∼ 19.0 (field 246); Sept. 8.89, ∼ 18.0 (field 298). This star is not present on the PSS prints (1953 Dec. 31) or on an OCA film obtained in 1994 Sept. (limiting mag 22).

Pollas also reports that SNe 1989H (cf. IAUC 4774) and 1989J (IAUC 4777) are both visible on the Palomar Sky Survey R plate (field 383), taken on 1989 Mar. 9.8 UT, at R ∼ 20.7 and 19.0, respectively.

COMET 122P/1995 S1 (DE VICO)

Extended ephemeris from orbital elements on MPC 25715:

1995TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Oct. 30 14^{h}12.^{m}97 +31^{o}28.^{′}7 1.124 0.820 45.^{o}1 59.^{o}0  6.9
Nov. 4 14 49.55 +30 32.7 1.210 0.881 45.9 53.9  7.4
9 15 20.21 +29 08.3 1.305 0.947 46.2 49.1  7.8
14 15 45.81 +27 33.3 1.406 1.015 46.2 44.7  8.3
19 16 07.32 +25 58.2 1.509 1.086 45.9 40.8  8.8
24 16 25.64 +24 28.5 1.612 1.158 45.4 37.4  9.2
29 16 41.44 +23 06.9 1.714 1.230 44.9 34.4  9.6
Dec. 4 16 55.27 +21 54.4 1.813 1.303 44.2 31.8  9.9
9 17 07.51 +20 50.9 1.909 1.376 43.6 29.6 10.3
14 17 18.48 +19 56.3 2.001 1.448 43.1 27.7 10.6
19 17 28.39 +19 10.1 2.088 1.520 42.8 26.1 10.9
24 17 37.41 +18 31.7 2.170 1.592 42.6 24.7 11.2
29 17 45.66 +18 00.6 2.247 1.663 42.7 23.6 11.5

1995 October 19 Daniel W. E. Green